A metal (M) forms a compound with the formula MI2 . The compound contains 79.97% I by mass. Identify the metal in the compound? Give the symbol of the metal.
1
Expert's answer
2019-10-25T05:30:59-0400

M(MI2)=x+127·2=x+254

W(I)=127·2/(x+254)=0.7997

x=63.6. Cu